# Comment Traps: How Defective Commented-out Code Augment Defects in AI-Assisted Code Generation

### v1

@@ Abstract.
This paper has been accepted by FSE 2026 (ACM International Conference on the Foundations of Software Engineering). This is a preprint version and may differ from the final published version.
$c$s: Software and its engineering Automatic programmingccs: Software and its engineering Software defect analysis
@@ 1. Introduction
The integration of large language models (LLMs) with Integrated Development Environments (IDEs) marks a transformative advancement in modern software engineering. Tools such as GitHub Copilot (b1), Cursor (b2), and related systems offer developers real-time, intelligent code suggestions, significantly enhancing coding efficiency. Empirical studies indicate that these tools can contribute up to 43 additional lines of code per session (b3), underscoring their effectiveness in collaborative programming contexts.
@@
Nonetheless, despite their substantial utility, recent research highlights notable security concerns in the code generated by such tools. A study of GitHub Copilot’s recommendations revealed that 39.33% of top suggestions and 40.73% of all suggestions contained vulnerabilities classified under MITRE’s Common Weakness Enumerations (CWEs) (b4). Subsequent evaluations show modest improvements in later versions, with the rate of vulnerable Python code suggestions declining from 36.54% to 27.25% (b6).
@@
However, prior research on code defects has mainly examined executable code (b4; b6; b44), neglecting the role of non-executable commented-out code (CO code), a factor that AI coding assistants may interpret as part of the prompt, thereby affecting the generated code generation. CO code refers to code segments temporarily disabled using language-specific comment syntax, allowing developers to retain nonexecutable code within source files (b7). It is commonly used for debugging, experimenting with alternative implementations, or preserving legacy logic for potential future use.
@@
Despite its developmental utility, the impact of CO code on AI-powered programming assistants remains underexplored. Since CO code often forms part of the prompt given to AI coding assistants, defects in this code may lead to the generation of additional flaws. We refer to such instances as defective CO code. While recent studies have examined the effects of natural language comments (e.g., TODOs) on code generation performance (b8), as well as the influence of contextual elements on GitHub Copilot’s effectiveness (b9; b10; b11), the role of CO code has not been systematically analyzed.
@@
We use a case study to illustrate the impact of defective CO code. Figure 1(a) shows a secure SSL connection code generated by GitHub Copilot. Lines 12 to 15, highlighted with a green background, contain content generated by GitHub Copilot based on the provided code prompt, which explicitly specifies a secure SSL version. Conversely, Fig. 1(b) reveals that inserting defective CO code in the prompt above the completion point (lines 11 to 14) leads GitHub Copilot to generate flawed code in lines 16 to 19. In this flawed segment, GitHub Copilot invokes the function ssl.wrap_socket without specifying the protocol version, potentially resulting in the use of deprecated or insecure TLS versions. This issue is identified as CWE-327 in the CWEs.
@@
Figure 1. A Case Study on the Impact of Defective CO Code on AI Coding Assistants
A Case Study on the Impact of Defective CO Code on AI Coding Assistants.
@@
To verify the prevalence of CO code in real-world development, we analyzed 6,403 high-quality Python repositories from GitHub, examining their CO code counts and identifying defects in CO code using CodeQL (b39). We found that 84.62% of repositories and 13.09% of Python files contained CO code, with 21.19% of those CO code having detectable defects. This highlights that a notable number of such issues exist in practice.
@@
Our main experimental procedure is illustrated in Fig. 2. We scanned Python files to identify defective code segments, which were then extracted and commented to create defective CO code. The remaining parts of the files were preserved as code context, forming a targeted dataset. Next, we inserted the defective CO code into different positions within the corresponding code context to generate prompts containing the defective CO code. These prompts were submitted to the AI coding assistants GitHub Copilot and Cursor, with code generation performed using the GPT-4o and Claude 3.5 Sonnet models. We then examined whether the generated code reintroduced the original defect. The results show that when defective CO code is included in the prompts, the number of defects in the generated code increases. We also evaluated the impact of factors such as tagging, truncation strategies, and the sparsity of the surrounding code context. Given that prompt engineering can influence the output of AI coding assistants (b45), we assessed its effectiveness in mitigating this issue. To guide this study, we formulated four research questions:
@@
RQ1: Impact of CO code on defect rates in AI-generated code. We first examined how defective CO code affects the number of defects in code generated by GitHub Copilot and Cursor. After inserting defective CO code into the code context, the number of defects generated by AI coding assistants exceeds that of the control group without inserting CO code and the control group with random CO code.
@@
RQ2: Impact of CO code truncation and tagging on defect rates in AI-generated code. Building on RQ1, we modified defective CO code to explore how AI tools interpret them. Removing 50% of trailing characters shows that tools still generate defects. Adding “<<Vulnerable>>” tags before or after the defective CO code has little effect.
@@
RQ3: Effectiveness of prompt engineering in mitigating CO code effects. Subsequently, we examined whether explicitly instructing the AI programming tool to avoid referencing defects mentioned in the defective CO code was effective. The practice of prompting the model to ignore defective CO code and regenerate yielded limited success, as only a small fraction of the outputs became defect-free.
@@
RQ4: Impact of code context sparsity on defect rates in AI-generated code influenced by CO code. Finally, we examined whether AI coding assistants’ parsing of inserted defective CO code is influenced by surrounding blank lines and indentation. The results indicate that for the GitHub Copilot tool, sparser context, such as code surrounded by blank lines, is associated with higher defect rates.
@@ 2.1. GitHub Repository Collection
To ensure the selection of recent, high-quality GitHub repositories, we applied the following filtering criteria: (1) Stars greater than 1,000, indicates broad community recognition and repository reliability. (2) The programming language is Python, we selected Python as the subject of our study due to its widespread use. (3) Size between 10 KB and 500,000 KB, excluding trivial or excessively large repositories that may contain non-source artifacts. (4) With commits after December 2022, ensure timeliness by collecting as many new code samples as possible to reduce the risk of data leakage in subsequent experiments. The final dataset includes 6,403 repositories, totaling 451 GB with 1,015,407 Python files and 648,016 folders.
@@ 3.1. AI-Assisted Code Generation
Large language models (LLMs) have achieved notable progress in natural language understanding and generation, demonstrating strong performance across diverse tasks (b12; b13; b14). They have also shown substantial capability in software engineering tasks such as code generation (b15), bug fixing (b17) and code completion (b18). Integrated AI coding assistants like GitHub Copilot enhance these capabilities through IDE integration, offering improved contextual awareness and productivity over standalone models or APIs (b21). Cursor is also a popular AI-powered programming tool, widely favored by world-class developers for its intelligent and efficient coding experience (cursor_features). Donvir et al. (10741797) noted that, in terms of generating secure code, Cursor offers enhanced security through context-aware code generation, while GitHub Copilot provides basic security suggestions.
However, recent studies highlight security risks in AI-generated code, a systematic survey conducted by Dwight Horne (b22) identifies multiple potential risks associated with the use of AI coding assistants. The study by Gavina Baralla et al. (b23) reveals limitations of GitHub Copilot in handling complex blockchain-specific logic and security considerations. Deniz Aydın et al. (11130176) investigated a range of security vulnerabilities present in JavaScript code generated by six LLMs, encompassing 28 distinct CWEs, an empirical study by Yujia Fu et al. (10.1145/3716848) on code generated by three AI code-generation tools identified 43 distinct CWEs. The study by Pearce et al. (b4), along with its replication study (b6), revealed that code generated by GitHub Copilot contains numerous defects detectable using CodeQL. These studies sound an alarm for AI code-generation tools, reminding us that while we harness their powerful code-generation capabilities, we must pay special attention to their ability to generate secure code.
@@ 3.2. Factors Influencing the Performance of AI-Generated Code
Several studies show that AI-generated code performance is influenced by multiple factors, Arora et al. (b25) investigated how different combinations of hyperparameters influence the performance of LLMs. Geng et al. (b26) introduced a semantic-guided pruning method that optimizes input length for LLMs, enhancing performance on the ClassEval benchmark. Al-Khafaji et al. (b27) assessed LLMs using Arabic prompts, revealing variations in code generation performance across models and prompt structures. Research also indicates that multi-step Chain-of-Descriptions improves the quality of VHDL code generation (b28), while ensembling expert models enhances Verilog code quality (b29). Fine-tuning is widely used to improve code generation in open-source models (b30; b32), whereas in-context prompting offers an efficient alternative for closed-source models (b34; b35). Heo et al. (b36) examined how the scope and format of contextual information affect code generation. They found that a broader context and the presence of comments improve generation quality. The study by Gavin S. Black et al. (10658990) investigated various types of prompts and found that appropriately incorporating safety guidance can effectively reduce vulnerabilities in generated code, with model selection exerting the most significant impact.
@@ 4. Methodology
Fig. 2 outlines the methodological workflow. When creating methodology flowcharts, we employed icons from open-access icon repositories, including icons for GitHub Copilot, and Cursor, among others (b40; b41; b42). The study proceeds through four phases:
@@ 4.1. Dataset Preparation
To make the dataset more representative of real-world scenarios, we built it using high-quality repositories collected from GitHub. We first used CodeQL (b39) to scan all Python files in these repositories for defects. Each file containing a defect became an entry in the dataset.
@@
This dataset classifies defects using CodeQL labels and metadata. These include 207 vulnerabilities, 48 reliability issues, 376 defects, 340 maintainability issues, 9 correctness issues, and 20 modularity issues. Vulnerabilities refer to defects classified under the Research Concepts view (CWE-1000) (cwe1000v415), covering 17 distinct CWE types belonging to 7 Pillars. We have annotated each vulnerability type in the dataset. This broad range of defect categories enhances the reliability and comprehensiveness of our study.
@@ 5. Experiments
Notably, when GPT-4o and Claude 3.5 Sonnet are used within Cursor, no significant LPAE is observed. However, LPAE consistently appears in GitHub Copilot regardless of the model used. This indicates that the effect is driven primarily by GitHub Copilot’s internal processing logic rather than the choice of LLM.
To understand this phenomenon, we examined GitHub Copilot’s official blog and documentation. Although the available information is limited and not fully up to date, we identified a likely contributing factor: the Fill-In-the-Middle (FIM) paradigm (rosenkilde2023how). Before FIM, only code preceding the cursor was included in the prompt, while subsequent code was ignored. FIM enables the model to distinguish between code before and after the cursor, allowing more accurate suggestions during non-linear editing. This advancement reportedly improved performance by 10%. However, by explicitly incorporating post-cursor context, FIM may increase the influence of defective code in later positions, thereby amplifying the introduction of defects.

### v2

@@ Abstract.
copyright: ccdoi: 10.1145/3797102journalyear: 2026journal: PACMSEjournalvolume: 3journalnumber: FSEarticle: FSE074publicationmonth: 7ccs: Software and its engineering Automatic programmingccs: Software and its engineering Software defect analysis
@@ 1. Introduction
The integration of large language models (LLMs) with Integrated Development Environments (IDEs) marks a transformative advancement in modern software engineering (Chen et al., 2025a; Huang et al., 2024). Tools such as GitHub Copilot (GitHub, 2021), Cursor (Cursor, 2023), and related systems offer developers real-time, intelligent code suggestions, significantly enhancing coding efficiency. Empirical studies indicate that these tools can contribute up to 43 additional lines of code per session (Imai, 2022), underscoring their effectiveness in collaborative programming contexts.
@@
Nonetheless, despite their substantial utility, recent research highlights notable security concerns in the code generated by such tools. A study of GitHub Copilot’s recommendations revealed that 39.33% of top suggestions and 40.73% of all suggestions contained vulnerabilities classified under MITRE’s Common Weakness Enumerations (CWEs) (Pearce et al., 2022). Subsequent evaluations show modest improvements in later versions, with the rate of vulnerable Python code suggestions declining from 36.54% to 27.25% (Majdinasab and others, 2024).
@@
However, prior research on code defects has mainly examined executable code (Pearce et al., 2022; Majdinasab and others, 2024; Li et al., 2024b; Huang et al., 2025; Chen et al., 2025b, 2024), neglecting the role of non-executable commented-out code (CO code), a factor that AI coding assistants may interpret as part of the prompt, thereby affecting the generated code generation. CO code refers to code segments temporarily disabled using language-specific comment syntax, allowing developers to retain nonexecutable code within source files (Pham and Yang, 2020). It is commonly used for debugging, experimenting with alternative implementations, or preserving legacy logic for potential future use.
@@
Despite its developmental utility, the impact of CO code on AI-powered programming assistants remains underexplored. Since CO code often forms part of the prompt given to AI coding assistants, defects in this code may lead to the generation of additional flaws. We refer to such instances as defective CO code. While recent studies have examined the effects of natural language comments (e.g., TODOs) on code generation performance (O’Brien et al., 2024), as well as the influence of contextual elements on GitHub Copilot’s effectiveness (Fagadau et al., 2024; Siddiq et al., 2023; Abukhalaf et al., 2023), the role of CO code has not been systematically analyzed.
@@
We use a motivating example to illustrate the impact of defective CO code. Figure 1(a) shows a secure SSL connection code generated by GitHub Copilot. Lines 12 to 15, highlighted with a green background, contain content generated by GitHub Copilot based on the provided code prompt, which explicitly specifies a secure SSL version. Conversely, Fig. 1(b) reveals that inserting defective CO code in the prompt above the completion point (lines 11 to 14) leads GitHub Copilot to generate flawed code in lines 16 to 19. In this flawed segment, GitHub Copilot invokes the function ssl.wrap_socket without specifying the protocol version, potentially resulting in the use of deprecated or insecure TLS versions. This issue is identified as CWE-327 in the CWEs.
@@
Figure 1. A Motivating Example on the Impact of Defective CO Code on AI Coding Assistants
A Motivating Example on the Impact of Defective CO Code on AI Coding Assistants.
@@
To verify the prevalence of CO code in real-world development, we analyzed 6,403 high-quality Python repositories from GitHub, examining their CO code counts and identifying defects in CO code using CodeQL (GitHub, 2025). We found that 84.62% of repositories and 13.09% of Python files contained CO code, with 21.19% of those CO code having detectable defects. This highlights that a notable number of such issues exist in practice.
@@
Our main experimental procedure is illustrated in Fig. 2. We scanned Python files to identify defective code segments, which were then extracted and commented to create defective CO code. The remaining parts of the files were preserved as code context, forming a targeted dataset. Next, we inserted the defective CO code into different positions within the corresponding code context to generate prompts containing the defective CO code. These prompts were submitted to the AI coding assistants GitHub Copilot and Cursor, with code generation performed using the GPT-4o (OpenAI, 2024) and Claude 3.5 Sonnet (Anthropic, 2024) models. We then examined whether the generated code reintroduced the original defect. The results show that when defective CO code is included in the prompts, the number of defects in the generated code increases. We also evaluated the impact of factors such as tagging, truncation strategies, and the sparsity of the surrounding code context. Given that prompt engineering can influence the output of AI coding assistants (Fagadau et al., 2024), we assessed its effectiveness in mitigating this issue. To guide this study, we formulated four research questions:
@@
RQ1: How does CO code impact defect rates in AI-generated code? We first examined how defective CO code affects the number of defects in code generated by GitHub Copilot and Cursor. After inserting defective CO code into the code context, the number of defects generated by AI coding assistants exceeds that of the control group without inserting CO code and the control group with random CO code.
@@
RQ2: How do truncation and tagging of CO code affect defect rates in AI-generated code? Building on RQ1, we modified defective CO code to explore how AI tools interpret them. Removing 50% of trailing characters shows that tools still generate defects. Adding “<<Vulnerable>>” tags before or after the defective CO code has little effect.
@@
RQ3: How effective is prompt engineering in mitigating the effects of CO code? Subsequently, we examined whether explicitly instructing the AI programming tool to avoid referencing defects mentioned in the defective CO code was effective. The practice of prompting the model to ignore defective CO code and regenerate yielded limited success, as only a small fraction of the outputs became defect-free.
@@
RQ4: How does code context sparsity influence defect rates in AI-generated code affected by CO code? Finally, we examined whether AI coding assistants’ parsing of inserted defective CO code is influenced by surrounding blank lines and indentation. The results indicate that for the GitHub Copilot tool, sparser context, such as code surrounded by blank lines, is associated with higher defect rates.
@@ 2.1. GitHub Repository Collection
To ensure the selection of recent, high-quality GitHub repositories, we applied the following filtering criteria: (1) Stars greater than 1,000, indicates broad community recognition and repository reliability. (2) The programming language is Python, we selected Python as the subject of our study due to its widespread use. (3) Size between 10 KB and 500,000 KB, excluding trivial or excessively large repositories that may contain non-source artifacts. (4) With commits after December 2022, ensure timeliness by collecting as many new code samples as possible. The final dataset includes 6,403 repositories, totaling 451 GB with 1,015,407 Python files and 648,016 folders.
@@ 3.1. AI-Assisted Code Generation
Large language models (LLMs) have achieved notable progress in natural language understanding and generation, demonstrating strong performance across diverse tasks (Abdullah et al., 2022; Rothman and Gulli, 2022; Tingiris and Kinsella, 2021). They have also shown substantial capability in software engineering tasks such as code generation (Qiuming et al., 2024), bug fixing (Bo et al., 2024) and code completion (Lai and others, 2024). Integrated AI coding assistants like GitHub Copilot enhance these capabilities through IDE integration, offering improved contextual awareness and productivity over standalone models or APIs (Corso et al., 2024). Cursor is also a popular AI-powered programming tool, widely favored by world-class developers for its intelligent and efficient coding experience (Cursor, 2025). Donvir et al. (Donvir et al., 2024) noted that, in terms of generating secure code, Cursor offers enhanced security through context-aware code generation, while GitHub Copilot provides basic security suggestions.
However, recent studies highlight security risks in AI-generated code, a systematic survey conducted by Horne (Horne, 2023) identifies multiple potential risks associated with the use of AI coding assistants. The study by Baralla et al. (Baralla et al., 2024) reveals limitations of GitHub Copilot in handling complex blockchain-specific logic and security considerations. Aydın et al. (Aydın and Bahtiyar, 2025) investigated a range of security vulnerabilities present in JavaScript code generated by six LLMs, encompassing 28 distinct CWEs, an empirical study by Fu et al. (Fu et al., 2025) on code generated by three AI code-generation tools identified 43 distinct CWEs. The study by Pearce et al. (Pearce et al., 2022), along with its replication study (Majdinasab and others, 2024), revealed that code generated by GitHub Copilot contains numerous defects detectable using CodeQL. These studies sound an alarm for AI code-generation tools, reminding us that while we harness their powerful code-generation capabilities, we must pay special attention to their ability to generate secure code.
@@ 3.2. Factors Influencing the Performance of AI-Generated Code
Several studies show that AI-generated code performance is influenced by multiple factors, Arora et al. (Arora et al., 2024) investigated how different combinations of hyperparameters influence the performance of LLMs. Geng et al. (Geng et al., 2024) introduced a semantic-guided pruning method that optimizes input length for LLMs, enhancing performance on the ClassEval benchmark. Al-Khafaji et al. (Al-Khafaji and Majeed, 2024) assessed LLMs using Arabic prompts, revealing variations in code generation performance across models and prompt structures. Research also indicates that multi-step Chain-of-Descriptions improves the quality of VHDL code generation (Vijayaraghavan and others, 2024), while ensembling expert models enhances Verilog code quality (Nadimi and Zheng, 2024). Fine-tuning is widely used to improve code generation in open-source models (Juhász, 2024; Li et al., 2024a), whereas in-context prompting offers an efficient alternative for closed-source models (Joshua and Sreejith, 2025; Li et al., 2024c). Heo et al. (Heo and Lee, 2024) examined how the scope and format of contextual information affect code generation. They found that a broader context and the presence of comments improve generation quality. The study by Black et al. (Black et al., 2025) investigated various types of prompts and found that appropriately incorporating safety guidance can effectively reduce vulnerabilities in generated code, with model selection exerting the most significant impact.
@@ 4. Methodology
Fig. 2 outlines the methodological workflow. When creating methodology flowcharts, we employed icons from open-access icon repositories, including icons for GitHub Copilot, and Cursor, among others (Secret Studio and Freepik, 2025; LobeHub, 2025; SVGRepo, 2025). The study proceeds through four phases:
@@ 4.1. Dataset Preparation
To make the dataset more representative of real-world scenarios, we built it using high-quality repositories collected from GitHub. We first used CodeQL (GitHub, 2025) to scan all Python files in these repositories for defects. Each file containing a defect became an entry in the dataset. To mitigate the risk of data leakage, we first restricted our data to samples updated after December 2022 to prioritize recency. Second, during subsequent prompt construction process, we transformed the original data by inserting a CO code, ensuring that the resulting prompts differ from the code in the original repository.
@@
This dataset classifies defects using CodeQL labels and metadata. To enhance coverage, this paper focuses on defects, rather than being limited exclusively to security vulnerabilities. These include 207 vulnerabilities, 48 reliability issues, 376 defects, 340 maintainability issues, 9 correctness issues, and 20 modularity issues. Vulnerabilities refer to defects classified under the Research Concepts view (CWE-1000) (The MITRE Corporation, 2025), covering 17 distinct CWE types belonging to 7 Pillars. We have annotated each vulnerability type in the dataset. This broad range of defect categories enhances the reliability and comprehensiveness of our study.
@@ 5. Experiments
Notably, when GPT-4o and Claude 3.5 Sonnet are used within Cursor, no significant LPAE is observed. However, LPAE consistently appears in GitHub Copilot regardless of the model used.
To understand this phenomenon, we examined GitHub Copilot’s official blog and documentation. Although the available information is limited and not fully up to date, we identified a likely contributing factor: the Fill-In-the-Middle (FIM) paradigm (Rosenkilde, 2023). Before FIM, only code preceding the cursor was included in the prompt, while subsequent code was ignored. FIM enables the model to distinguish between code before and after the cursor, allowing more accurate suggestions during non-linear editing. This advancement reportedly improved performance by 10%. However, by explicitly incorporating post-cursor context, FIM may increase the influence of defective code in later positions, thereby amplifying the introduction of defects.
@@ 8. Data Availability
Acknowledgements.
The work described in this paper was supported by the National Key Research and Development Program of China (2023YFB2703700), the National Natural Science Foundation of China (62372492), and GMCC-SYSU Joint Lab for Smart Applications.